Construct an amortization schedule for a 3 year loan of 50000 at 5% interest , which is to be repaid in quaterly installment over 3 years.
1
Expert's answer
2021-09-06T16:05:22-0400

To get installments we use the following formula:

PV"=" PMT"(" "\frac {1-(1+r)^{-n}}{r})"

PV"=" 50000

r"=" interest rate"(" 0.05/4")""=" 0.0125

n is the time (3)

50000"=" PMT"(" "\frac {1-(1+0.0125)^{-3}}{0.0125}" ")"

PMT"=" 17085.06


To get interest for year 1

5000"\u00d7" 0.0125"=" 625

Principal"=" Installment "-" Interest

17085.06"-" 625"=" 16460.06

Outstanding balance"=" 50000"-"16460.06"=" 33539.94

For year 2 we will use first year's balance to calculate interest. Repeat the same for year 3